OpenStudy (anonymous):

Factor completely. 150v^2+360vd+216d^2

OpenStudy (anonymous):

\[150v^2+360vd+216d^2\] \[=6(25v^2+60vd+36d62)\] \[=6(5v+6d)^2\]